Solve the separable differential equation (y2+2)dx+y(x+4)dy=0.
Answer #1

- ztu (y²+2) dx + y (2+4) dy = y (x+4) dy = -(y?+2)+7 9 - da dy = tu y+z on integration ydy da f y2+2 Į Put y²+2=t gydy dt Yd

Thumbs up please

??